zunairah:

In 2014 a class went to the holocaust museum. They traveled in vans and busses. Each van held 10 students and each bus held 15 students. 14 vehicles total wwent on the trip. If 180 students went on the trip, how many of each type of vehicle did we use.

QuestionCoveBot:

So first off... you gave to divide what?

zunairah:

wwhat

QuestionCoveBot:

What do we divide here?

zunairah:

idk

QuestionCoveBot:

Okay... we divide 180 and 14. That gives your average in each vehicle. What do you get when you divide 180 and 14? 180/14

Rigo:

first how many buses went on the trip

Rigo:

then find out how many student swent total

AZ:

okay since no one here knows how to really solve the question lol, I'll take over The real way to solve this question is by coming up with two equations

AZ:

what letters or variables do you want to use? @zunairah x and y or V for vans and B for bus

AZ:

We make the two equations from the information they gave us Each van held 10 students and each bus held 15 students. 14 vehicles total went on the trip. 180 students went on the trip The first equation is based on 14 vehicles total went on trips that means number of vans + number of buses = 14 so depending on what variables/letters you want to use, you can replace it accordingly

AZ:

The silence is deafening but I will presume you are afk and continue The second equation would be using Each van held 10 students and each bus held 15 students. 180 students went on the trip How many students can a van hold? And a bus? 15 in a bus and 10 in a van and a total of 180 students went so 15*bus + 10*van = 180 Does that make sense?

AZ:

basically the variable is going to tell us how many of the buses we took or how many vans we took so when we multiply 15*bus it will tell us how many people took the bus in all and similarly for the vans

AZ:

and since you haven't chosen any variables, I'll just stick with V = # of vans B = # of buses so we have two equations V + B = 14 10V + 15B = 180 Do you know how to solve a system of equations? @zunairah

zunairah:

what

zunairah:

oh

zunairah:

yes i do

AZ:

did all of that make sense? how I came up with the equations? can you solve for V and B now?

zunairah:

ummm

AZ:

we can use x and y instead x = # of vans y = # of buses so we have two equations x + y = 14 10x + 15y = 180

zunairah:

yeah i did that and after that i got 11x+16y =194

AZ:

you can't just add the equations like that if you're doing substitution, you need to do something so that way one of the letters gets ELIMINATED

AZ:

so that way you can solve for just one variable at a time

zunairah:

so i multiply the top equation by something and the bottom equation by something

AZ:

yes, but in this case, you only have to multiply the first equation by something to get it to have the same number as the second one so you can cancel out one of the letters so do you want to get rid of x or y first?

zunairah:

x

AZ:

okay, so what would you need to multiply the first equation with? we have 10x in the second one we want the first one to start with -10x then so what do you multiply to get -10x from x

zunairah:

-10

AZ:

Good, so multiply the entire equation on both sides by -10 x + y = 14 -10 *(x+y) = -10 * 14 what do you get

zunairah:

-10x-10y=-140

AZ:

good now let's add the two equations -10x-10y=-140 10x + 15y = 180

zunairah:

i got 8 for y

AZ:

correct! and so if y is 8, then what is x? remember x + y = 14

zunairah:

6 for x

AZ:

and you got it!! and remember we said: x = # of vans y = # of buses so that means 6 vans and 8 buses

zunairah:

you remind me of my math teacher 😡 and thx bye

AZ:

LOL yw bye